# Pressure Drop in Non-Spherical Packed Beds: Influence of Geometry and Reynolds Number

### v2

@@ 5.5 Transition to Inertial Regime
Figure 12: Relative deviation of the friction factor from the $f_{\mathrm{p}}$. The circular markers with black outlines show the critical Reynolds number for each angle.
@@
Hlushkou2006 suggested that if the inertial contribution exceeds $5\%$, the flow should no longer be considered in the Darcy regime. The same criterion is adopted in the present study. To identify the Reynolds number at which inertial effects become significant, the difference between the total friction factor and the linear friction factor ($f_{\mathrm{lin}}=1/(\phi\,\mathrm{Re}_{\mathrm{p}}\,\mathrm{Da})$) is computed and then divided by the total friction factor. We applied the transition criterion proposed by Hlushkou2006 to identify the onset of inertia-dominated flow. However, other criteria have also been reported in the literature, such as introduced by Zeng2006; bagci2014. In future work, different transition criteria may be used to provide a more comprehensive assessment of the transition to inertia-dominated flow.
@@
At $\mathrm{Re}_{\mathrm{p}}=200$, the rotation angles $\alpha\in\{0^{\circ},5^{\circ},\dots,25^{\circ}\}$ and $\{85^{\circ},90^{\circ}\}$ deviate notably from the behaviour observed at the remaining angles. For these other angles, the nonlinear component accounts for 80–90% of the total friction factor. The primary mechanism driving this distinct behaviour is the formation of channel-like flow paths. Although some of these angles (such as $25\text{\,}\mathrm{\SIUnitSymbolDegree}$ and $85\text{\,}\mathrm{\SIUnitSymbolDegree}$) do not form fully developed channel-like geometries, they still generate localized preferential jet-like flow within the void space.
@@
At $\alpha=$55\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the transition to the inertial regime occurs at approximately $\mathrm{Re}_{\mathrm{p}}\approx 5$, which is lower than for the other geometries. Potential factors contributing to this behavior include the sensitivity of the Darcy number evaluation for this orientation and the unique connectivity of the void space, which may promote early localized recirculation. A conclusive explanation would require a detailed flow field analysis which currently exceeds the scope of this study.
@@ 6 Conclusion
In this study, numerical simulations were performed using OpenFOAM-12 to investigate viscosity- and inertia-dominated flow through periodically arranged square-bar structures over a range of $\mathrm{Re}_{\mathrm{p}}=0.1\text{--}200$ with different angles ($0\text{\,}\mathrm{\SIUnitSymbolDegree}$$90\text{\,}\mathrm{\SIUnitSymbolDegree}$) at fixed porosity $\phi=0.332$. A total of 266 simulations were performed using a predominated hexahedral mesh, and selected cases were successfully validated against available PIV measurements obtained from the experimental configuration of christin2025. Some important conclusions can be drawn from the obtained results:
@@
    The rotation angle has a strong influence on the macroscopic flow characteristics. For $\alpha\leq$20\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the structure forms channel-like geometries, resulting in high effective permeability over the entire $\mathrm{Re}_{\mathrm{p}}$ range and low flow resistance at high $\mathrm{Re}_{\mathrm{p}}$. As the angle increases, the connectivity between successive layers becomes more complex, leading to the development of recirculation zones and increased frictional losses, resulting in heterogeneous velocity fields. We refer to these configurations as lattice-like geometries. At $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the structure again forms a mainly channel-like arrangement. In this orientation, the flow path becomes a less perturbed, and the disturbances observed at other rotation angles are reduced.
@@
    For the investigated Reynolds numbers, the onset of te inertial regime was defined as the critical Reynolds number, which was identified for each angle by checking when the inertial contribution to the friction factor exceeded 5%. For lattice-like geometries, the transition occurs at $\mathrm{Re}_{\mathrm{p}}\approx 7.5$, with the earliest onset at $\alpha=$55\text{\,}\mathrm{\SIUnitSymbolDegree}$\$. When the rotation angle exceeds $20\text{\,}\mathrm{\SIUnitSymbolDegree}$, the inertial regime begins at higher Reynolds numbers.
@@
    At $\mathrm{Re}_{\mathrm{p}}=0.1$, the highest friction factor was obtained at $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, whereas at $\mathrm{Re}_{\mathrm{p}}=200$, it occurred at $\alpha=$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$. This demonstrates that the geometrical configuration exerts a strong influence on flow resistance, even at constant porosity.
@@
    It is important to note that the present numerical study was conducted at a fixed porosity to effectively highlight the impact of the inertial structure induced by the rotation angle. This influence of varying geometry is typically not included in standard permeability models, which are generally formulated as a function of porosity. Future work should consider modelling this geometric influence using approaches similar to those of (duplessis1988; VANWACHEM2024104815), validated against the present dataset.
Overall, the good agreement between simulations and experiments shows that the present configuration provides a reliable framework for analyzing flow through such periodic structures. Future work will extend the analysis to higher Reynolds numbers to examine the onset of unsteady and turbulent flow in more detail.
7 Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
8 Acknowledgments
The authors gratefully acknowledge the financial support of the Deutsche Forschungsgemeinschaft (DFG) through No. SFB/ TRR287, Project No. 422037413.
@@ Appendix A Effect of Mesh Resolution on the Numerical Results
Figure 13: Normalized spanwise-averaged velocity profiles at P3 (see fig. 1) for angles $30\text{\,}\mathrm{\SIUnitSymbolDegree}$, $60\text{\,}\mathrm{\SIUnitSymbolDegree}$, and $90\text{\,}\mathrm{\SIUnitSymbolDegree}$. Results are shown for three mesh resolutions (200K, 1M, and 4M cells)
Figure 14: Friction factor for three mesh sizes (200K, 1M, and 4M cells) at angles of $30^{\circ}$, $60^{\circ}$, and $90^{\circ}$.
Table 3: Friction factors used in for each angle and mesh. The table contains the number of cells per layer ($N_{\mathrm{C}}$)
@@
The analysis was conducted for the case shown in fig. 13, focusing on the first module of the computational domain along the probe line located at point P3 (see fig. 1), positioned at the mid-line in the $z$-direction. Velocity data were extracted from statistically steady-state solutions without applying periodic averaging. Figure 13 presents the normalized spanwise-averaged velocity profiles. The coarsest mesh ($2\text{\times}{10}^{5}$ cells) shows a noticeable deviation from the finer ones, particularly in the central region of the $30\text{\,}\mathrm{\SIUnitSymbolDegree}$ case. For $60\text{\,}\mathrm{\SIUnitSymbolDegree}$ and $90\text{\,}\mathrm{\SIUnitSymbolDegree}$, the deviation becomes evident near the peak locations of the velocity profiles, where the $2\text{\times}{10}^{5}$ cell mesh results in less accurate results. As the mesh is refined, the profiles converge toward those of the finest mesh, indicating improved resolution of the flow field. For all investigated angles, the $1\text{\times}{10}^{6}$ and $4\text{\times}{10}^{6}$ meshes yield nearly identical profiles, suggesting that the flow features are well-resolved from the $1\text{\times}{10}^{6}$ cell mesh.
@@
The mesh sensitivity analysis demonstrates that the $1\text{\times}{10}^{6}$ cell mesh provides reliable results while maintaining a reasonable computational cost.

### v3

@@ 5.6 Transition to inertial regime
Figure 15: Relative deviation of the friction factor from the Darcy-regime value. The circular markers with black outlines show the critical Reynolds number for each angle.
@@
Hlushkou2006 suggested that if the inertial contribution exceeds $5\%$, the flow should no longer be considered in the Darcy regime. The same criterion is adopted in the present study. To identify the Reynolds number at which inertial effects become significant, the difference between the total friction factor and the linear friction factor ($f_{\mathrm{lin}}=1/(\phi\,\mathrm{Re}_{\mathrm{p}}\,\mathrm{Da})$) is computed and then divided by the total friction factor. However, other criteria have also been reported in the literature, such as introduced by Zeng2006; bagci2014. In future work, different transition criteria may be used to provide a more comprehensive assessment of the transition to inertia-dominated flow.
@@
[…]
At $\alpha=$55\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the transition to the inertial regime occurs at approximately $\mathrm{Re}_{\mathrm{p}}\approx 5$, which is lower than for the other geometries. Potential factors contributing to this behaviour include the sensitivity of the Darcy number evaluation for this orientation and the unique connectivity of the void space, which may promote early localized recirculation. A conclusive explanation would require a detailed flow field analysis which currently exceeds the scope of this study.
@@ 6 Conclusion
In this study, numerical simulations were performed using OpenFOAM-12 to investigate viscosity- and inertia-dominated flow through periodically arranged square-bar structures over a range of $\mathrm{Re}_{\mathrm{p}}=0.1\text{--}200$ with different angles ($0\text{\,}\mathrm{\SIUnitSymbolDegree}$$90\text{\,}\mathrm{\SIUnitSymbolDegree}$) at fixed porosity $\phi=0.332$. A total of 266 simulations were performed using a predominantly hexahedral mesh, and selected cases were successfully validated against available PIV measurements obtained from the experimental configuration of christin2025. Some important conclusions can be drawn from the obtained results:
@@
    The rotation angle has a strong influence on the macroscopic flow characteristics. For $\alpha\leq$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the structure forms channel-like geometries, resulting in high permeability. As the angle increases beyond $10^{\circ}$, the connectivity between successive layers becomes more complex, leading to the development of recirculation zones and increased frictional losses, resulting in heterogeneous velocity fields. We refer to these configurations ($\alpha\geq 15^{\circ}$, including $90^{\circ}$) as lattice-like geometries. The $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ configuration is also classified as lattice-like, as confirmed by its Darcy number, bed-to-particle diameter ratio and tortuosity.
@@
    Using a 5% inertial contribution threshold, the transition to the inertial regime occurs at $\mathrm{Re}_{\mathrm{p}}\approx 7.5$ for lattice-like geometries, with the earliest onset at $\alpha=$55\text{\,}\mathrm{\SIUnitSymbolDegree}$\$.
@@
    The drag characteristics are strongly impacted by the geometric configuration, and the angle yielding the highest friction factor shifts with the Reynolds number. In the viscous regime, the severe constriction of the flow path at $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ results in the highest friction factor. In the inertial regime, the contraction–expansion sequences at $\alpha=$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ lead to flow separation and recirculation, resulting in the highest friction factor.
@@
    The classification into channel-like and lattice-like regimes can be supported quantitatively by the bed-to-particle diameter ratio and the hydraulic tortuosity. Geometries with $D/d_{\mathrm{eq}}<3$ and $\tau<1.1$ behave as channel-like, whereas those with $D/d_{\mathrm{eq}}>3$ and $\tau>1.1$ fall into the lattice-like group.
  5. (5)
    The module-equivalent diameter $d_{\mathrm{eq}}$, which accounts for the angle-dependent wetted surface area, collapses the friction factor data onto the Ergun correlation more effectively than the constant single-bar diameter $d_{\mathrm{sb}}$, allowing to partially account for the geometry variation. Among the permeability models, the Blake–Kozeny estimate based on $d_{\mathrm{eq}}$ provides the best agreement with the DNS data for the lattice-like geometries (MAPE $12.3\%$). The tortuosity-based estimate $K_{\tau}$ (Kozeny constant $c_{\mathrm{KC}}=3.1$) reproduces the angle-dependent trend but with larger deviations (MAPE $18.9\%$), while the single-bar estimate yields substantially larger errors (MAPE $80.2\%$). All three models underestimate the permeability of the channel-like configurations ($\alpha\leq 10^{\circ}$), as they are formulated for multiply connected pore structures.
  6. (6)
    The Forchheimer coefficient $C_{\mathrm{F}}$ varies strongly with rotation angle, peaking at $\alpha=60^{\circ}$ and remaining low for the channel-like and $90^{\circ}$ configurations. At intermediate angles ($30^{\circ}$$70^{\circ}$), $C_{\mathrm{F}}$ fluctuates most strongly, reflecting the sensitivity of inertial losses to the void-space topology in this range.
The good agreement between simulations and experiments confirms that the present approach provides a reliable framework for analysing flow through such periodic structures. Future work should extend the analysis to higher Reynolds numbers to examine the onset of unsteady and turbulent flow, and include additional configurations to improve the fitting of geometry-aware models for both permeability and the Forchheimer coefficient.
Acknowledgments
The authors gratefully acknowledge the financial support of the Deutsche Forschungsgemeinschaft (DFG) through No. SFB/ TRR287, Project No. 422037413. The authors wish to thank Christin Velten, Kerstin Hülz, and Katharina Zähringer for sharing the dataset of the experimental measurements of flow in the packed bed.
@@ Appendix A Effect of mesh resolution on the numerical results
Figure A.1: Normalized spanwise-averaged velocity profiles at P3 (see Fig. 1) for angles $30\text{\,}\mathrm{\SIUnitSymbolDegree}$, $60\text{\,}\mathrm{\SIUnitSymbolDegree}$, and $90\text{\,}\mathrm{\SIUnitSymbolDegree}$. Results are shown for three mesh resolutions (200K, 1M, and 4M cells)
Figure A.2: Friction factor for three mesh sizes (200K, 1M, and 4M cells) at angles of $30^{\circ}$, $60^{\circ}$, and $90^{\circ}$.
Table A.1: Friction factors used in for each angle and mesh. The table contains the number of cells per layer ($N_{\mathrm{C}}$)
@@
The analysis was conducted for the case shown in Fig. A.1, focusing on the first module of the computational domain along the probe line located at point P3 (see Fig. 1), positioned at the mid-line in the $z$-direction. Velocity data were extracted from statistically steady-state solutions without applying periodic averaging. Figure A.1 presents the normalized spanwise-averaged velocity profiles. The coarsest mesh ($2\text{\times}{10}^{5}$ cells) shows a noticeable deviation from the finer ones, particularly in the central region of the $30\text{\,}\mathrm{\SIUnitSymbolDegree}$ case. For $60\text{\,}\mathrm{\SIUnitSymbolDegree}$ and $90\text{\,}\mathrm{\SIUnitSymbolDegree}$, the deviation becomes evident near the peak locations of the velocity profiles, where the $2\text{\times}{10}^{5}$ cell mesh results in less accurate results. As the mesh is refined, the profiles converge toward those of the finest mesh, indicating improved resolution of the flow field. For all investigated angles, the $1\text{\times}{10}^{6}$ and $4\text{\times}{10}^{6}$ meshes yield nearly identical profiles, suggesting that the flow features are well-resolved from the $1\text{\times}{10}^{6}$ cell mesh.